A widow currently has a $100,000 investment that yields 5 percent annually. Can she withdraw $24,000 for the next five years? Use Appendix D to answer the question. Round your answer to the nearest dollar. The maximum amount that can be withdrawn is $ so she withdraw $24,000 for the next five years. Would your answer be different if the yield were 8 percent? Use Appendix D to answer the question. Round your answer to the nearest dollar. If the yield is 8 percent the maximum amount that can be withdrawn is $ so she withdraw $24,000 for the next five years.
Answer #1

1

PVOrdinary Annuity = C*[(1-(1+i/100)^(-n))/(i/100)]
C = Cash flow per period
i = interest rate
n = number of years
PV= 24000*((1-(1+ 5/100)^-5)/(5/100))
PV = 103907
Using Calculator: press buttons "2ND"+"FV" then assign
PMT =24000
I/Y =5
N =5
FV = 0
CPT PV
Using Excel
=PV(rate,nper,pmt,FV,type)
=PV(5/(100),5,,PV,)

No as PV is more than 100000

2

PVOrdinary Annuity = C*[(1-(1+i/100)^(-n))/(i/100)]
C = Cash flow per period
i = interest rate
n = number of years
100000= Cash Flow*((1-(1+ 5/100)^-5)/(5/100))
Cash Flow = 23098 = max amount
Using Calculator: press buttons "2ND"+"FV" then assign
PV =-100000
I/Y =5
N =5
FV = 0
CPT PMT
Using Excel
=PMT(rate,nper,pv,fv,type)
=PMT(5/(100),5,,100000,)

3

PVOrdinary Annuity = C*[(1-(1+i/100)^(-n))/(i/100)]
C = Cash flow per period
i = interest rate
n = number of years
PV= 24000*((1-(1+ 8/100)^-5)/(8/100))
PV = 95825
Using Calculator: press buttons "2ND"+"FV" then assign
PMT =24000
I/Y =8
N =5
FV = 0
CPT PV
Using Excel
=PV(rate,nper,pmt,FV,type)
=PV(8/(100),5,,PV,)

Yes as PV is less than 100000

4

PVOrdinary Annuity = C*[(1-(1+i/100)^(-n))/(i/100)]
C = Cash flow per period
i = interest rate
n = number of years
100000= Cash Flow*((1-(1+ 8/100)^-5)/(8/100))
Cash Flow = 25046 = max cash flow
Using Calculator: press buttons "2ND"+"FV" then assign
PV =-100000
I/Y =8
N =5
FV = 0
CPT PMT
Using Excel
=PMT(rate,nper,pv,fv,type)
=PMT(8/(100),5,,100000,)